Unit 1 Review:

1) Find inverse of g(x) = 3x+1/ 2

1
Expert's answer
2021-09-20T18:23:26-0400
"g(x)=3x+\dfrac{1}{2}"

Replace "g(x)" and "y"


"y=3x+\dfrac{1}{2}"

Change "x" and "y"


"x=3y+\dfrac{1}{2}"

Solve for "y"


"y=\dfrac{1}{3}x-\dfrac{1}{6}"

Replace "y" and "g^{-1}(x)"


"g^{-1}(x)=\dfrac{1}{3}x-\dfrac{1}{6}"